# LongVILA: Scaling Long-Context Visual Language Models for Long Videos

### v3

@@ 2 GPUs 4 GPUs 8 GPUs
Figure 10: Comparison of Needle in the Long Video Haystack Experiment. The 32-frame baseline model (left) can not retrieve right needles after 32 frames. In contrast, the LongVILA model (right), trained on 1024 frames, presents 99.5% accuracy on 274k context length.
@@ 5.1.2 Inference system
We evaluated our inference system against HuggingFace Pipeline parallelism using a single node with 8 H100 GPUs and the 8B model (Table  3 and Figure 9 (b)). Our system achieves an 8.2×8.2$\times$8.2 × speedup compared to HuggingFace Pipeline on 8xH100 GPUs. This significant improvement is primarily due to HuggingFace Pipeline inference activating only one GPU at a time, whereas our method leverages all GPUs to compute jointly. Table 3 compares the maximum supported sequence length, where our method supports sequences that are 2.9×2.9$\times$2.9 × longer than those supported by HuggingFace Pipeline. Specifically, during 96K sequence length inference, the maximum sequence length supported, HuggingFace Pipeline stores 80GB of activations on the first GPU and only 18GB on the remaining GPUs. This imbalanced allocation of activations limits the maximum sequence length that can be processed.
@@ 5.1.3 Effect of two-stage sharding
We evaluate the impact of our two-stage sharding strategy using a video captioning dataset (Chen et al., 2024a). We compare our two-stage sharding approach to a baseline one-stage sharding strategy that only distributes workload based on the number of images. We measure the time per iteration across different numbers of H100 GPUs. For $k$ GPUs, we use $k$ images per video and a batch size of $k$. The results, shown in Table 4, indicate a speedup ranging from 1% to 7%. This improvement is primarily observed in longer captioning tasks, where the baseline suffers from workload imbalance due to the lack of sharding based on the number of text tokens.
Table 5: Performance evaluation on Video MME (Fu et al., 2024a) benchmark.
Model Frames Short Medium Long Overall
Qwen-VL-Chat 4 46.9 38.7 37.8 41.1
Video-LLaVA (Lin et al., 2023a) 8 45.3 38.0 36.2 39.9
ST-LLM (Liu et al., 2024b) 64 45.7 36.8 31.3 37.9
VideoChat2-Mistral (Li et al., 2023b) 16 48.3 37.0 33.2 39.5
Chat-UniVi-V1.5 (Jin et al., 2023) 64 45.7 40.3 35.8 40.6
VideoLLaMA2 (Cheng et al., 2024) 16 56.0 45.4 42.1 47.9
LLaVA-NeXT-Qwen2 32 58.0 47.0 43.4 49.5
LongVILA-8B 128 60.2 48.2 38.8 49.2
256 61.8 49.7 39.7 50.5
Table 6: Ablation on Stage 3: Short SFT and Stage 4: Mid-training for LongVILA-8B on Video MME benchmark with 128 frames for evaluation.
Stage 3 Stage 4 Average Short Medium Long
@@
Table 7: Evaluation of LongVILA-Caption performance across different frame counts. LongVILA-Caption’s performance improves across all metrics. Unlike in Figure 2, the models presented in this table are both trained and evaluated using the same number of frames.
@@ 5.2 Needle-in-a-Haystack
5.3 General Video Understanding
Table 5 presents the performance of various models on the Video MME benchmark, comparing their effectiveness across short, medium, and long video lengths, as well as overall performance. LongVILA-8B, utilizing 256 frames, achieves an overall score of 50.5. We also do the ablation on the effects of Stage 3 and Stage 4 in Table 6.
@@ 6 Conclusion
We introduce LongVILA, a comprehensive full-stack solution for long-context visual language models, encompassing model training pipeline and distributed system. Based on our curated long video datasets and five-stage training pipeline, our LongVILA-8B model extends the feasible frame count from 8 to 1024, precisely capturing fine-grained information from 2-hour needle-in-a-haystack videos and delivering promising results in both video QA and captioning tasks. Additionally, our system efficiently scales context length up to 2 million tokens, achieving speedups of 2.1$\times$ to 5.7$\times$ compared to ring sequence parallelism and 1.1$\times$ to 1.4$\times$ compared to Megatron context parallelism + tensor parallelism.
@@ Appendix A Appendix
Figure 12: Convergence evaluation. We compare the training loss curves for LongVILA with and without sequence parallelism. This figure illustrates the convergence of the LongVILA-8B model on 8 A100 GPUs, with a sequence parallelism degree of 4, compared to pure data parallelism training. The training dataset includes a mixture of shot2story, ai2d, and chartqa. The two curves align closely, indicating that our MM-SP system does not negatively impact training quality.
@@
FSDP versus Zero-3 with hybrid sequence parallelism

### v4

@@ 5.1.2 Inference system
We evaluated our inference system against HuggingFace Pipeline parallelism using a single node with 8 H100 GPUs and the 8B model (Figure 10). Our system achieves an 8.2×8.2$\times$8.2 × speedup compared to HuggingFace Pipeline on 8xH100 GPUs. This significant improvement is primarily due to HuggingFace Pipeline inference activating only one GPU at a time, whereas our method leverages all GPUs to compute jointly. Figure 10 compares the maximum supported sequence length, where our method supports sequences that are 2.9×2.9$\times$2.9 × longer than those supported by HuggingFace Pipeline. Specifically, during 96K sequence length inference, HuggingFace Pipeline stores 80GB of activations on the first GPU and only 18GB on the remaining GPUs. This imbalanced allocation of activations limits the maximum supported sequence length.
@@ 5.1.3 Effect of two-stage sharding
We evaluate the impact of our two-stage sharding strategy using a video captioning dataset (Chen et al., 2024a). We compare our two-stage sharding to a one-stage baseline that only distributes workload based on the number of images. We measure the time per iteration across different numbers of H100 GPUs. For $k$ GPUs, we use $k$ images per video and a batch size of $k$. The results, shown in Table 2, indicate a speedup ranging from 1% to 7%. This improvement is primarily observed in longer captioning tasks, where the baseline suffers from workload imbalance due to the lack of sharding based on the number of text tokens.
Table 3: Performance comparison on VideoMME (Fu et al., 2024a) benchmark.
Model LLM Size Frames w/o subtitle w subtitle
Overall Short Medium Long Overall Short Medium Long
Video-LLaVA 7B 8 39.9 45.3 38.0 36.2 41.6 46.1 40.7 38.1
SliME 8B 8 45.3 53.3 55.4 39.8 47.2 55.4 44.4 41.7
ShareGPT4Video 8B 16 39.9 48.3 36.3 35.0 43.6 53.6 39.3 37.9
VideoChat2 7B 16 39.5 48.3 37.0 33.2 43.8 52.8 39.4 39.2
VideoLLaMA2 7B 16 47.9 56.0 45.4 42.1 50.3 59.4 47.6 43.8
Chat-Univi-v1.5 7B 64 40.6 45.7 40.3 35.8 45.9 51.2 44.6 41.8
Kangaroo 8B 64 56.0 66.1 55.3 46.7 57.6 68.0 55.4 49.3
ShareGemini 7B 64 43.2 49.1 41.3 39.1 47.9 49.1 47.3 43.4
LongVA 7B 128 52.6 61.1 50.4 46.2 54.3 61.1 53.6 47.6
InternVL-V1.5 20B 10 50.7 60.2 46.4 45.6 52.4 61.7 49.1 46.6
VITA 8x7B 20 55.0 64.2 53.3 47.6 57.6 67.9 55.3 49.6
Video-CCAM 14B 96 53.9 62.1 52.8 47.0 56.1 63.9 55.9 48.3
LongVILA 7B 128 57.5 69.3 56.1 47.0 60.6 70.4 59.2 52.1
256 57.4 69.0 56.7 46.4 61.8 70.8 60.6 54.0
Table 4: Ablation on Stage 3 (S-3) and Stage 4 (S-4) for LongVILA on VideoMME.
S-3 S-4 Average Short Medium Long
@@
Table 5: Evaluation of LongVILA-Caption performance across different frame counts.
@@
5.2 General Video Understanding
Table 3 presents the performance of LongVILA, comparing to state-of-the-art models (Lin et al., 2023a; Zhang et al., 2024c; Chen et al., 2024a; Li et al., 2023c; Cheng et al., 2024; Jin et al., 2023; Zhang et al., ; Share, 2024; Chen et al., 2024c; Fu et al., 2024b; Fei et al., 2024; Liu et al., 2024c) on the VideoMME benchmark, comparing their effectiveness across short, medium, and long video lengths, as well as overall performance. LongVILA, utilizing 128 frames, achieves an overall score of 57.5 / 60.6 without / with subtitle, which are competitive results. We also do the ablation on the effects of Stage 3 and Stage 4 in Table 5, using Llama3 based VILA models.
@@ 6 Conclusion
We introduce LongVILA, a comprehensive full-stack solution for long-context visual language models, encompassing model training pipeline and distributed system. Based on our curated long video datasets and five-stage training pipeline, our LongVILA model extends the feasible frame count from 8 to 2048, precisely capturing fine-grained information from 2-hour needle-in-a-haystack videos and delivering promising results in both video QA and captioning tasks. Our system efficiently scales context length up to 2 million tokens, achieving speedups of 2.1$\times$ to 5.7$\times$ compared to ring sequence parallelism and 1.1$\times$ to 1.4$\times$ compared to a hybrid Megatron context and tensor parallelism.
@@ Appendix A Appendix
Figure 12: Convergence evaluation. We compare the training loss curves for LongVILA with and without sequence parallelism. This figure illustrates the convergence of the LongVILA model on 8 H100 GPUs, with a sequence parallelism degree of 4, compared to pure data parallelism training. 2U+2Z indicates that 2D-Attention mechanism is enabled, where both the Ulysses and Zigzag-Ringttn degrees are 2. The training dataset is shot2story. The two curves align closely, indicating that our MM-SP system does not negatively impact training quality.
@@
Table 8: Comparison with state-of-the-art methods Li et al. (2023b); Dai et al. (2023); Bai et al. (2023); Liu et al. (2023b); Lin et al. (2023b); Liu et al. (2024b); Tong et al. (2024); Li et al. (2024) on 10 image based VLM benchmarks. S3 refers to the stage 3 model in LongVILA training pipeline.
[…]
BLIP-2 Vicuna-13B 224 41.0 41 19.6 61 42.5 46.4 38.1 22.4
InstructBLIP Vicuna-7B 224 49.2 34.5 60.5 50.1 36 23.7 53.4 60.9 26.2
Vicuna-13B 224 49.5 33.4 63.1 50.7 58.2 25.6
Qwen-VL Qwen-7B 448 78.8 59.3 35.2 67.1 63.8 38.2 7.4 56.3
Qwen-VL-Chat Qwen-7B 448 78.2 57.5 38.9 68.2 61.5 60.6 56.7 58.2
LLaVA-1.5 Vicuna-1.5-7B 336 78.5 62.0 50.0 66.8 58.2 64.3 58.3 58.6 63.4 30.5
Vicuna-1.5-13B 336 80.0 63.3 53.6 71.6 61.3 67.7 63.6 61.6 70.7 35.4
VILA Llama 2-7B 336 79.9 62.3 57.8 68.2 64.4 68.9 61.7 61.1 69.7 34.9
Llama 2-13B 336 80.8 63.3 60.6 73.7 66.6 70.3 64.3 62.8 73.0 38.8
LLaVA-NeXT-8B Llama 3-8B 672 65.2 72.8 64.6 72.1 80.1
Cambrian-1-8B Llama 3-8B 1024 64.6 80.4 71.7 75.9
Mini-Gemini-HD-8B Llama 3-8B 1536 64.5 75.1 70.2 72.7
LongVILA-7B (S3) Qwen2-7B dynamic 85.4 65.4 65.0 98.5 77.8 83.4 80.0 70.6 77.6 51.7